OpenStudy (anonymous):

find the slope of the line tangent to the curve r=2cos3θ, at θ=π/6

OpenStudy (campbell_st):

can you find the 1st derivative...?

OpenStudy (anonymous):

-6sin3θ

OpenStudy (campbell_st):

great... the 1st derivative is the equation of the slope of the tangent... so substitute the value \[\theta = \frac{\pi}{6}\] this will give the slope...

OpenStudy (campbell_st):

what value did you get for the slope..?

OpenStudy (anonymous):

-6 ?

OpenStudy (campbell_st):

correct... thats all you need... remember the 1st derivative gives the equation of the slope of the tangent. for a specific slope you need to substitute the x value from the point.

OpenStudy (anonymous):

thanks